SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Sept 6, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Fang Fang, Larry Gladney, Selina Li, Dong Su

Operations

  • no major trigger problem until this morning when the power supply for the BLT/PTD crate failed. See more info in the DCT section.

EMT

  • The EMT prescale test failed again last Wednesday during the MD. See the ROM logfile here. Rainer said he could help out with this. Dan needs to check in the code so that we can look at the code to see what can be done to resolve the problem.
  • The EMT test stand is in working condition since last Thursday so that Dan can use it to test loading the prescale configuration at the test stand.

DCT/DCZ

  • At 6:12am this morning, we got paged about the DCT crate not recognizing links (and "All values whited out.") See the ROM logfile here. We suggested the pilot to take DCT out of partition and take data again. No errors seen other than the expected display in old vs new TSF comparison fastmon plots. From run 57801 to 57805, the DCT was out of partition in physics data-taking. The pilot included the DCT in configuration again after the power supply was replaced by Ray.
  • A few fastmon problems were found for those runs taken without the old DCT:
    • Some of the fastmon codes need to be independent of the old DCT system. An example is the GLT input lines plot which checks for PTD digits. This needs to be fixed.
    • The page for comparison plots for BLT was broken (in page 9 of the fastmon archive plots).
    • This one is not related to the DCT, but it's in the archive plots. The time history plot for event rate almost always has the first bin out of scale. Rainer suggested to throw away the first incomplete bin.
    Basically, we need to switch off the fastmon codes that monitor the old DCT data.
  • We decided to remove the old DCT system from data-taking sometime this week. Other than creating a crate mask excluded the DCT, we also should make sure that the fastmon processors don't expect to see the old DCT data (thus creating error messages). We will also make sure that the old DCT can be turned back on again in November.

GLT

  • Su Dong is not sure the GLT-FC firmware now uses 10 or 20 microsecond DAQ delay. This needs to be tested at the test stand to find out how big the gap is. He is still somewhat surprised the GLT still contribute quite a bit deadtime.

L3

  • No more node trims.

Data Quality and Monitoring

  • The time history plots were included in fastmon archive since last Wednesday. It couldn't be implemented at JAS due to the scalar display problem. Rainer suggested to ask online event processing about how to implement it.
  • For fastmon, Su Dong pointed out that the cable check shouldn't check the damaged events. This will clean up the disagreement seen in the GLT-FCT corruptions.
  • L3 plots for rates, lumi, etc will be made automatically by Jim Hamilton soon (still on his to-do-list).
  • SP8 validation still needs to be done yet. Waiting for Eric (who is probably on vacation now) to sign off.

Trigger configuration

  • Two PHYSICS_L1PASSTHRU runs (with off-resonance data) 57689 and 57690 were taken last week for trigger rate studies. We will actually need to take the runs with on-resonance data. However, this is not possible until later this week when we are running on-resonance again and also until the configuration problem was fixed. Apparently, the only person who knows how to fix this configuration database problem is Andy who is on vacation now. Rainer also pointed out that the second run could be a background run rather than a real physics run due to the mess up in loading new configuration.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.